OpenStudy (anonymous):

Find the equation of the tangent to the curve y = e^x and that is parallel to the line x - 4y = 1

OpenStudy (anonymous):

First you must have to find the derivative of the curve. rewrite the other equation to determine the slope -4y = -x+1 y = 1/4x -1 so m = 1/4 do you know how to find the derivative of the curve?